-Resubmitted due to typo.
Need help please. This question makes no sense, and I can't find an example that explains what needs to be done to solve it.
n(U) = 100, n(A) = 24, n(B) = 60, and n(A intersection B) = 10. Find n(A' intersection B).
Thank you!

Found 2 solutions by itxmenaved, ikleyn:
Answer by itxmenaved(3)

n(A')=n(u)-n(A)=100-24
n(A')=76
now from N(AUB)=10 these 10 elements are not is A'
(because an elemet is in A is not Not In A')
and the commont elemts in A' and B are 50
so n(A' intersection B)=50

Answer by ikleyn(52803)

.
It was solved and carefully explained under this link

https://www.algebra.com/algebra/homework/sets-and-operations/sets-and-operations.faq.question.1129419.html

https://www.algebra.com/algebra/homework/sets-and-operations/sets-and-operations.faq.question.1129419.html